### v10

# Power Set Axiom, Internal Classes, External Sets

Abstract
Usual math sets have special types: countable, compact, open, occasionally Borel, rarely projective, etc. Generic sets dependent on Power Set axiom appear mostly in esoteric areas,
logic of Set Theory (ST), etc. Recognizing internal to math (formula-specified) and external
(based on parameters in those formulas) aspects of math objects greatly simplifies foundations.
I postulate external sets (not internally specified, treated as the domain of variables) to be hereditarily countable and independent of formula-defined classes, i.e. with finite Kolmogorov Information about them. This allows elimination of all non-integer quantifiers in ST formulas.
ZFC axioms are sometimes given to undergraduates in an unintuitive,
hard to remember list. Setting them in three pairs seems to help intuition.
I always wondered why math foundations as taken by logicians are so distant from the actual math practice. For instance, the cardinality theory - the heart of the Set Theory (ST) - is almost never used beyond figuring out which sets are countable and which are not.
I see the culprit in the Power Set axiom (somewhat similar to “set of all sets”, albeit not so fatal). Usual math objects are sets of special types: countable, compact, open, occasionally Borel, rarely projective, etc.00footnotemark: 0 They by nature are not really sets but classes, defined by ST formulas with countable sets as parameters. Collections of them are really collections of those parameters.
I attribute many complications in Math foundations to blurred distinction between two types of collections different in nature. One are pure classes defined by math formulas. The other are external objects that math handles but generally does not specify internally, for instance, random sequences. Externals are proper sets and are the domain of ST variables. I postulate them to be hereditarily countable and independent of pure classes, i.e. have only finite Kolmogorov Information about them. Math objects are mixed classes specified by math formulas with external sets as parameters.
Excluded would be “Logics” objects such as, e.g., the collection of all true statements of Arithmetic.
ST power can ever be expanded with stronger axioms. There is no natural end to this endeavor.
A more feasible goal is a tighter match to the actual math usage. Studying “higher level” theories
is exciting but can be done without claiming relation to the foundations of mainstream math.
1 A Smaller Purge: Against Pathologies
Many reversals of Power Set axiom to “all sets are countable” (resulting in some forms of second order arithmetic) have been studied so far (e.g., Kripke-Platek ST, pocket ST [2], reverse math [2, 6], etc.). Yet even in that part some aspects here need to differ, due to a different goal. It is not in basing some specific areas on weaker assumptions, with different, possibly more complicated, proofs. Rather it is to fit most of mainstream math, with little, if any, substantive changes in proofs. Quantifiers here will bind all sets (no cases restricted to $\bm{\Delta}_{0}$, for instance) but not all classes. Classes (only informally called sets) are meta-objects, not technically in the domain of the variables.
Formalities.
This, in effect, splits math objects into a hierarchy of ranks, each quantifier binding only classes below one rank. Rank ${\leq}k$ classes $F^{k}_{p}$ are specified by a hereditarily countable (“proper” set) parameter $p$ via the universal $\Sigma^{1}_{k}$ formula $F^{k}_{p}(q)$. In $p$, an ordinal $o(p)$ is included. $F^{k}_{p}(q)$ defines membership $F^{k}_{q}{\in}F^{k}_{p}$ conditioned on $o(q){<}o(p)$. Membership is also extended by extensionality: $F^{k^{\prime}}_{q^{\prime}}{\in}F^{k}_{p}$ if $F^{k}_{q}$, $F^{k^{\prime}}_{q^{\prime}}$ have the same members, so have a formula-defined isomorphism of membership relations on their transitive closures. Typical math concepts have straightforward translations.
For instance, “$x$ is in open $P{\subset}{\mathbb{R}}$” can be a shorthand for “$F^{k}_{p}(x)$, where $p$ specifies $P{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}F^{k}_{p}$ as
the set of its rational intervals”. (Or $P$ can be Borel, or whatever type clear from the context.)
2 A More Radical Purge: Against Tricksome Objects
And making explicit the internal to math (defined by formulas) and external (parameters based) aspects of math objects clarifies their nature. Even for countable collections this allows a better focus on issues brought by these two different (albeit with many similarities) sources.
“Pure” classes defined by parameter-free ST formulas are specific but tricksome.
Careless treatment of them easily brings paradoxes, but careful treatment complicates the theory. They should not be combined in further definable collections as that would only extend the language of allowed formulas. A formula can take parameters for external data (sets) which math handles without specifying internally. Sets constitute the domain of ST variables.
Sets can be put in collections based on properties and relations with other sets. This forms “mixed” classes, specified by ST formulas with some free variables taken for data parameters. Data may be chaotic, but simpleminded, playing no self-referential tricks. General math objects – mixed classes – carry both tricks and chaos. But their tricks are limited to what a single formula can carry.
In reality, external data end up finite. Yet handling their (often ambiguous) termination points is awkward. So, countably infinite data are convenient. [^1] An example of more technical conveniences: the ring (class) of p-adic integers, is compact: “less infinite” than ${\mathbb{Z}}$. But distinguishing them (sets) from pure classes allows another radical insight: Even with infinite complexity, external objects $\alpha{\in}\,{\mathbf{\Omega}}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{0,1\}^{\mathbb{N}}$ can have only finite (small, really) dependence on pure sequences $\Psi$, i.e. Kolmogorov Information ${\mathbf{I}}(\alpha:\Psi)<\infty$. (Sec. 3 gives some background.) Note, recursive and random transformations can only add finite information about a given (e.g., by a math formula) sequence.
See [4] for discussion in its section “The Independence Postulate” (IP) and for references.
Due to this Postulate (IP), only computable sequences can double as both pure classes and sets.
IP conflicts with Specification axioms for sets. (For classes these axioms are merely definitions.) So, they must be restricted to just one: “Being a set is preserved by Turing reductions”. Instead,
the Foundation axiom becomes a family, like Induction axioms: one for each class of ordinals.
[…]
Generic are $\gamma{\in}\,{\mathbf{\Omega}}$ outside all arithmetic sets $X{\in}\Delta^{0}_{*}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\cup_{n}\Delta^{0}_{n}$ of measure $\lambda(X){=}0$. Take one. The model includes all $\alpha$ computable by algorithms $A$ from $\gamma^{(k)}$ where $\gamma^{(k)}_{i}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}0$ if $k|i$, else $\gamma^{(k)}_{i}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\gamma_{i}$. (In this model, 2nd order quantifiers are easy to eliminate, proving $P\Leftrightarrow\lambda(\{\gamma{:}P^{\prime}(\gamma)\}){>}0$ where $P^{\prime}$ has all real variables $\alpha_{i}$ in a closed $P$ replaced with $A_{i}(\gamma^{(k_{i})})$, and $A_{i},k_{i}$ quantified as integers. Cf.[3].)
So, mixed math objects would not encroach on infinite hierarchies but be reducible to a pair of a pure string (defined by a single formula) and a string carrying no information about pure ones.
3 Kolmogorov Information: A Brief Background
Uniform measure on ${\mathbf{\Omega}}$ (or on ${\mathbf{\Omega}}^{2}{\simeq}{\mathbf{\Omega}}$) is $\lambda(x\,{\mathbf{\Omega}}){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}2^{-\|x\|}$. (For $x{\in}\{0,1\}^{n}$, $\|x\|{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}n$; $x{\sqsubset}\alpha{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\alpha{\in}x{\mathbf{\Omega}}$.)
Partial continuous transforms (PCT) on ${\mathbf{\Omega}}$ may fail to narrow-down the output to
a single sequence, leaving a compact set of eligible results. So, their graphs are compact sets
$A\subset{\mathbf{\Omega}}{\times}{\mathbf{\Omega}}$ with $A(\alpha){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{\beta:(\alpha,\beta){\in}A\}\neq\emptyset$. Singleton outputs $\{\beta\}$ are interpreted as $\beta{\in}\,{\mathbf{\Omega}}$.
Preimages $A^{-1}(s){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{\alpha:A(\alpha)\subset s\}$ of all open sets $s\subset{\mathbf{\Omega}}$ in any PCT $A$ are open.
Closed $A$ also have closed preimages of all closed $s$. For $x{\in}\{0,1\}^{*}$ let $x^{A}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{\alpha:A(\alpha)\in x{\mathbf{\Omega}}\}$.
Computable PCTs have algorithms enumerating the clopen subsets of ${\mathbf{\Omega}}^{2}{\setminus}A$. $U(p\alpha)$ is a universal PCT. It computes $n{=}\|x\|$ bits of $A_{p}(\alpha){\subset}x{\mathbf{\Omega}}$ in $t_{p\alpha}(n)$ steps (and $U(p){\in}{\mathbb{N}}$ in $t_{p}$ steps).
A computably enumerable (c.e.) function to $\overline{{\mathbb{R}}^{+}}$ is $\sup$ of a c.e. set of basic continuous ones.
Dominant in a convex class $C$ of functions is its c.e. $f\in C$ if all c.e. $g$ in $C$ are $O(f)$.
Such is $\sum_{i}g_{i}/(i^{2}{+}i)$ if $(g_{i})$ is a c.e. family of all c.e. functions in weakly compact $C$.
$\lambda$-test is ${\mathbf{d}}(\alpha){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\|\lceil{\mathbf{T}}(\alpha)\rceil\|$ for a dominant ${\mathbf{T}}$ on ${\mathbf{\Omega}}$ with expectation $\lambda({\mathbf{T}}){\leq}1$. $\|t\|{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\lceil\log_{2}t\rceil{-}1$.
Martin-Lof $\lambda$-random are $\alpha{\in}{\mathbf{R}^{\lambda}}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}{\mathbf{R}^{\lambda}_{\infty}}$, where ${\mathbf{R}^{\lambda}_{c}}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{\alpha:{\mathbf{d}}(\alpha){<}c\}$ (compact for $c{\in}{\mathbb{N}}$).
Let ${\mathbf{M}}(x){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\lambda(U^{-1}(x{\mathbf{\Omega}}))$. ${\mathbf{R}^{\lambda}}$ consists of all $\gamma$ with $\sup_{x\sqsubset\gamma}\frac{{\mathbf{M}}(x)}{\lambda(x\,{\mathbf{\Omega}})}<\infty$.
Mutual Information (dependence) ${\mathbf{I}}(\alpha_{1}{:}\alpha_{2})$ is $\min_{\beta_{1},\beta_{2}}\{{\mathbf{d}}(\beta_{1},\beta_{2}):U(\beta_{i}){=}\alpha_{i}\}$.
Measuring classes.
For a sequence of clopen $F_{i}{\subset}{\mathbf{\Omega}}$ with ${\mathbf{M}}(F_{i}{\oplus}F_{i+j}){<}2^{-i}$, let open $F^{+}_{i}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\cup_{j>i}F_{j}$,
compact $F^{-}_{i}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\cap_{j>i}F_{j}$; $F^{-}\!{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\cup_{j>i}F^{-}_{j}{\subset}F^{+}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\cap_{j>i}F^{+}_{j}\!$. Any $F{\in}\Delta^{0}_{*}$ has such $(F_{i}){\in}\Delta^{0}_{*}$ with $F^{-}{\subset}F{\subset}F^{+}\!$.
Note that no generic $\gamma\in F^{+}\setminus F^{-}$. If $\lambda(F)=0$ then, for $F\subset{\mathbf{R}^{\lambda}}$, also ${\mathbf{M}}(F_{i})<2^{-i}$, implying ${\mathbf{I}}(\gamma:(F_{i})_{i}){=}\infty$ for $\gamma{\in}F\cap{\mathbf{R}^{\lambda}}$. Thus, any arithmetic class $F$ includes no generic $\gamma{\in}{\mathbf{R}^{\lambda}}$, if $\lambda(F){=}0$, and includes all such $\gamma$ or none if $F$ is invariant under single digit flips.
3.1 Reductions to Generic Sequences Turned Weak Truth-table (Closed PCT)
Let $s_{t}^{n}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\lambda(\{\alpha:t_{\alpha}(n){<}t\})$, $s_{\infty}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\inf_{n}s^{n}_{\infty}$. The function $\tau_{r}^{n}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\min\{t:s_{t}^{n}{>}$ $r{\in}{\mathbb{Q}}\}$ is computable. Note that if ${\mathbf{I}}(\alpha{:}\,s_{\infty}){<}\infty$ then $s_{\alpha}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\liminf_{n}s^{n}_{t_{\alpha}(n)}{<}s_{\infty}$ as $s_{\infty}{\notin}\Delta^{0}_{2}$. Take $r{\in}(s_{\alpha},s_{\infty})$ and a monotone infinite sequence $n^{\alpha}_{i}$ of all $n$ with $t_{\alpha}(n){<}\tau_{r}^{n}$. If ${\mathbf{I}}(\alpha{:}\,s_{\infty}){<}\infty$ then $\forall^{\infty}i\;\min\{p:t_{p}{>}n^{\alpha}_{i}\}<i\|i\|^{2}$.
$U_{c}(\alpha){\in}\{\#,0,1\}^{\mathbb{N}}$ avoids divergence by diluting $U(\alpha)$ with $\min_{p<i\|i\|^{2}+c}\,\{t_{p}{:}\;t_{p}{>}n^{\alpha}_{i+1}\}{\leq}\infty$ blanks ##\## after $n^{\alpha}_{i}$-th bits. $U_{c}$ carries no extra information of $\alpha$ absent in $U(\alpha)$, except only for $n^{\alpha}$.
As $U_{c}$ never diverges, $\mu{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}U_{c}(\lambda)$ is computable and $U^{\prime}(\alpha){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\mu([\#^{\mathbb{N}},U_{c}(\alpha)])$ maps ${\mathbf{R}^{\lambda}}$ to ${\mathbf{R}^{\lambda}}\cup U({\mathbb{N}})$.
From $\beta=U^{\prime}(\alpha){\in}{\mathbf{R}^{\lambda}}$ we recover $U_{c}(\alpha)$ and $u(\beta){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}U(\alpha)$. PCT $u$ is closed (a w.t.t. reduction)
as the input segments it uses are only slightly longer (by codes $p$ for $n^{\alpha}$ bounds) than the output’s.
(Viewing $\#^{+}$ segments of $U_{c}(\alpha)$ as commas makes $\mu$ a computable (for ${\mathbb{N}}^{*}$ intervals) distribution on sequences of integers. $U^{\prime}$ gives short codes to these integers #+$\#^{+}\{0,1\}^{+}$ and their prefixes.)
4 Elimination of 2nd Order Quantifiers with IP
[…]
integer quantifiers, we need to reduce $\{\beta:\exists\alpha\,(\alpha,\beta){\in}P\}$, for any reduced $P$, where $\beta\in{\mathbf{\Omega}}^{k}{\simeq}{\mathbf{\Omega}}$.
Note, $\exists\alpha{\in}{\mathbf{\Omega}}\,P(\alpha)\Leftrightarrow\exists\alpha{\in}{\mathbf{R}^{\lambda}_{0}}\,P(u(\alpha))$, so $\alpha,\beta,\gamma$ below will range only over ${\mathbf{R}^{\lambda}}$. Also, any pair $(\alpha,\beta)$ is $(u(\gamma,0),u(\gamma))$, so we need to reduce $u(P^{\prime})$ for reduced $P^{\prime}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\{\gamma{:}\;(u(\gamma,0),u(\gamma)){\in}P\}$.
Let $\lambda_{x}(Q){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\frac{\lambda(Q\cap x^{u})}{\lambda(x^{u})},$ $\lambda_{\beta}(Q){\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\liminf_{x\sqsubset\beta}\lambda_{x}(Q)$ be $\lambda$ on $\gamma$ conditioned on $u(\gamma){=}\beta$. Note: $\lambda_{\beta}(Q){>}t$ is reduced if $Q$ is. IP assures $\lambda_{\beta}(P^{\prime}){>}0$ for $\beta{\in}u(P^{\prime})$, so taking $P_{k}{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}$ $P^{\prime}\cup\{\gamma{:}\;\lambda_{u(\gamma)}(P^{\prime}){\leq}2^{-k}\}$ turns $\beta{\in}u(P^{\prime})$ into $\exists k\,(\beta{\in}u(P_{k})\&\,\lambda_{\beta}(P^{\prime}){>}2^{-k})$. For $F{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}P_{k}$, $\forall x\exists i\,\lambda_{x}(F^{-}_{i}){>}2^{-k}$. We are to reduce $u(F^{-})$ or $u(F^{-}_{i}){\cap}{\mathbf{R}^{\lambda}_{i}}$. A straight way to show ${\mathbf{R}^{\lambda}}{\subset}u(F^{-})$ is proving that ${\mathbf{R}^{\lambda}_{0}}$ has no one-way functions:
Recursively one-way functions.
Let $\lambda(X){>}0$, $A(\alpha){\in}{\mathbf{R}^{\lambda}_{0}}$ for all $\alpha{\in}X$. A closed computable PCT $A$ is one-way on $X$ if no computable $B$ can invert it, i.e. if $\lambda^{2}(\{(\beta,\gamma):\alpha{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}B(\beta,\gamma){\in}X,A(\alpha){=}\beta\})=0$. The question is, do such one-way functions exist? (Considering compact $X\in\Pi^{0}_{3}$ suffices.)
A substitute
can be extending an axiom. By IP, $(\exists\alpha\,\alpha{\in}X{\subset}{\mathbf{R}^{\lambda}})\Rightarrow\lambda(X){>}0$ for pure $X{\in}\Delta^{0}_{*}$.
[…]
5 Some Discussion
Early self-referential paradoxes showed problems with collections like “all true formulas” in domains covered by quantifiers. [^3] A subclass of formulas is safer, but seems an artificial complication of doubtful necessity. The concept of sets may open a backdoor if problematic collections of formulas can be included. Balancing such issues may be the source of various problems in math foundations based on Set Theory. One may be Cardinality Theory, almost never used in math beyond telling apart countable and uncountable sets. And it seems not very real as all theories have countable models. It might be just a way to obscure self-referential nature of some sets descriptions.
A clean way out may be recognizing the distinction between internal to math collections specified by its formulas and external ones that math needs to handle without specifying. Internal collections have a limited hierarchy: the type of allowed formulas is clear-cut. Any extension would make a new theory, with its own clear limits. External objects would be fully independent of internal ones: Formally, infinite Kolmogorov Information between the two types is proscribed.
Generic objects can be mixed: collections specified by formulas using external sets as parameters. So, any such object would rely on only one formula, not on all of them, thus provably excluding a back-door for extended language of formulas. Of course, such treatment would require well thought out conventions to minimize the effects on regular math practice.444 I am not clear yet on how to handle some math theorems that depend on Axiom of Choice, as it provides classes not specified by ZF formulas. One option may be extending the language with a postulated (not otherwise described) class that maps v.Neumann ordinals (may be just countable ones, implying Continuum Hypothesis) onto reals.
But any such options would require careful analysis.
But I do not see why this would require a very different care level than proper use of ZFC now does.
Acknowledgments.
I am grateful to Albert Meyer for helpful criticism.
Appendix: ZFC Axioms
Sets with a given ST property $P$ (possibly with parameters $c$) are said to form a class $\{x:P_{c}(x)\}$. They may or may not form a set, but only sets are the domain of ZFC variables.
  1. 1.
    Chains axioms.
    1a. Infinity. A non-empty set exists included in the union of its elements,
    i.e. having no sinks in its membership graph  [^5] Infinite ordinals are built by unique ordinal-valued function $f(s)=\sup_{x\in s\cap S}(f(x){+}1)$ on $S$. : $\exists S,s{\in}S\forall x{\in}S\exists y{\in}S(x{\in}y)$.
    1b. Foundation/Induction (anti-dual to 1a) : $\neg\exists S,s{\in}S\forall x{\in}S\exists y{\in}S(y{\in}x)$.
    Each non-empty set has membership graph sources (members disjoint with it).
  2. 2.
    Formula-defined sets.
    2a. Extensionality. Assigning members (as in 2b) defines sets uniquely: sets with the same members are members of the same sets : $\forall t(t{\in}x\Leftrightarrow t{\in}y)\Rightarrow\forall t(x{\in}t\Rightarrow y{\in}t)$.
    So, $x{=}y$ for sets or classes denotes $x{\subset}y{\subset}x$. ($x{\subset}y$, or $y{\supset}x$ mean $\forall t(t{\in}x\Rightarrow t{\in}y)$.)
    2b. Replacement family (an axiom for each ST-defined relation $R_{c}(x,y)$).
    The image $R_{c}(X)\;{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\;\{y:\exists x{\in}X\,R_{c}(x,y)\}$ of any set $X$ is a set if
[…]
    (or $R_{x}(x,y)$ if not using Powerset) plus Paring axiom. (Build $X{\times}\{c\}$. Conversely, $2^{a},2^{b},\{a,\{\}\},\{a,b\}$.)
    : $(\forall x\exists Y\supset R_{c}(\{x\}))\Rightarrow\forall X\exists Y{=}R_{c}(X)$.
  3. 3.
    Inverses of functions. Define functions $f$, their domains $\delta_{f}$, and classes $f^{-1}$ of their
    inverses $g:f(\delta_{f})\to\delta_{f}$, such that $f(g(x))=x$ (so, $g\subset f^{T}\;{\raisebox{-2.0pt}{$\,\stackrel{{\scriptstyle\mbox{\tiny df}}}{{=}}\,$}}\;\{(f(x),x):x{\in}\delta_{f}\}$).
    3a. Choice. All functions have inverses : $\forall f\exists\,g\in f^{-1}$.
    (The feasibility of computing inverses is the most dramatic open CS problem.)
    3b. Powerset. Assures these huge classes $f^{-1}$ of inverses $g\subset f^{T}$ still are sets :
    $\forall x\exists Y\forall z(z{\subset}x\Rightarrow z{\in}Y)$.
The modifications discussed above include:
[…]
(2) Turning Foundation and Extensionality into families of axiom: one for each class,
(3) restricting Replacement to recursive $R_{c}(X)$, (4) reversing Powerset to “all sets are countable”, and (5) some (not clear to me yet) replacement for Choice: possibly as discussed in 00footnotemark: 0.